# Quasiparticle cooling algorithms for quantum many-body state preparation

### v1

@@ D.2 $\sigma^{\pm}_{j;\alpha;\beta}$ matrix elements
We now turn to quasiparticle processes mediated by the $\hat{\sigma}^{\pm}_{j}$ matrix elements. We focus on $\hat{\sigma}^{-}_{j}$ for illustration. In the fermion basis, $\hat{\sigma}^{-}_{j}$ comes attached with the Jordan-Wigner string operator,
$\sigma^{-}_{j;\vec{\alpha};\vec{\beta}}=\bra{\vec{\alpha}}\Big{(}\prod_{i=1}^{j-1}e^{i\pi\hat{c}^{\dagger}_{i}\hat{c}_{i}}\Big{)}\hat{c}^{\dagger}_{j}\ket{\vec{\beta}}.$ (94)
Due to this non-locality in the fermionic eigenbasis, multi-particle processes are allowed with $\hat{\sigma}_{j}^{-}$ changing the occupation numbers of up to $12j-$ 1 quasiparticle levels. Amazingly, the exact finite-volume matrix elements have been worked out for the case of the $\hat{X}_{j}$ operators in the TFIM with periodic boundary conditions [77]. Here, we focus on singe- and two-particle transitions near the ground state, which capture the dominant cooling processes at late times, as described in the main text. Our theory can be straightforwardly generalised to include higher-order processes.
@@
$\sigma^{-}_{j;0;k}=\bra{\Omega}\hat{\sigma}^{-}_{j}\ket{k}$ (98)
being the amplitude to directly excite a quasiparticle of quasimomentum $k$. We rewrite the single-particle element in fermion basis
$|\sigma^{-}_{j;0;k}|^{2}=|\langle\prod_{i=1}^{j-1}\hat{a}_{2i-1}\hat{a}_{2i}\ \hat{c}^{\dagger}_{j}\hat{\eta}^{\dagger}_{k}\rangle_{\Omega}|^{2}.$ (99)
@@
Figure 10: (a) Single-quasiparticle matrix elements $|\sigma^{-}_{j0k}|^{2}$ summed over quasimomenta $k$ and plotted against operator position $j$, for different values of $J/g$. (b) Two-particle matrix elements $|\sigma^{-}_{j0(k,q)}|^{2}$ summed over $k$, $q$, and plotted against operator position $j$. Note that the two-particle matrix element is non-zero only in the AFM phase.
Two-quasiparticle processes: Based on the physical picture, we expect that in the AFM phase, a pair of domain walls may collide in the bulk and annihilate, as this requires only local reconfigurations of spins. However, there is a subtlety, from the fact that $\hat{\sigma}^{-}_{j}$ has non-zero matrix elements only between eigenstates of different fermion parity. This rules out two-particle transitions in the PM phase, but as discussed in the main text, the issue is sidestepped in the AFM by interacting with the Majorana zero mode, whose action maps between the two nearly-degenerate ground states. We therefore write
$|\sigma^{-}_{j;0;k,q}|^{2}=|\bra{\Omega_{+}}\prod_{i=1}^{j-1}\hat{a}_{2i-1}\hat{a}_{2i}\ \hat{c}^{\dagger}_{j}\hat{\eta}^{\dagger}_{k}\hat{\eta}^{\dagger}_{q}\ket{\Omega_{-}}|^{2},$ (100)
with |$\ket{\Omega_{-}}=\hat{\eta}^{\dagger}_{0}\ket{\Omega_{-}}$, and $\hat{\eta}^{\dagger}_{0}$ the Majorana mode. These matrix elements may be evaluated with Wick’s theorem on the $(2j+2)\times(2j+2)$ dimensional $\Sigma$-matrix. In Fig. 10b, we show how the summed matrix elements $\sum_{k,q}|\sigma^{-}_{j;0;k,q}|^{2}$ depend on the position $j$, for different values of $J/g$ in the AFM phase. The pair matrix element is almost homogeneous in the bulk, vanishing in the edge region.
@@
$|\sigma^{-}_{j;k;q}|^{2}=|\bra{\Omega_{+}}\hat{\eta}_{k}\prod_{i=1}^{j-1}\hat{a}_{2i-1}\hat{a}_{2i}\ \hat{c}^{\dagger}_{j}\hat{\eta}^{\dagger}_{q}\ket{\Omega_{-}}|^{2}.$ (101)
@@ E.1 Free fermion simulations
The reset of the auxiliaries can be done by reinstating condition (105). If we denote the corresponding map by $\mathcal{R}$, then the cooling channel (1) acts on the covariance matrix according to (in a slight abuse of notation):
@@ E.2 Tensor network simulations
Here, we summarize the tensor network algorithm that was employed to simulate the dissipative dynamics in Sections III-V. We give an overview of the relevant details, assuming the reader is familiar with the basics of matrix product algorithms; introductions can be found in e.g. [80, 81].
@@
The time evolution is performed via the standard TEBD algorithm [64, 65], which updates the MPS in real time, followed by the reset channel acting on the bath, which can be efficiently implemented as discussed below. In Fig. 11 we illustrate the evolution for the example of the Floquet TFIM (5) with bath evolution and interaction, Eqs. (7-10). We first simultaneously perform the single-site unitary channels for the system and bath, then the two-body system channels, which correspond to next-nearest-neighbor gates, followed by the nearest-neighbor system-bath channel. Note that the next-nearest-neighbour gate arises from the quasi-1d geometry. This is the most expensive step of the algorithm as it requires two singular value decompositions and scales as $O(\text{dim}(m)^{4}d^{3})$.
@@
Next, we study the effects of truncation due to finite bond dimension during the cooling protocol in Figures 12 and 13. In Fig. 12 we consider the TFIM cooling for the parameters shown in Fig. 3 of the main text. We only show data for two values of the coupling, $\theta=0.1,0.05$ for the MCP, since the SCP protocol is well converged for all bond dimensions. We observe that the convergence is slower for the larger coupling, suggesting higher bond dimensions are needed to capture non-perturbative effects. Phenomenologically, such behavior is expected since higher-order in $\theta$ corrections are expected to lead processes that are less spatially local, thus requiring a larger bond dimension to be accurately captured.
@@
Appendix F Comparison of noisy kinetic equation and MPS simulations
In Section IV we derived the kinetic equation in the presence of noise, Eq. (51), which includes transitions induced by qubit dephasing and decay, Eqs. (51). Here we compare the accuracy of the rate equation predictions against MPS simulations. We focus on the MCP with the same parameters as in Fig. (4), namely finite density of auxiliaries, $N_{A}=N_{S}=20$, and $T=12=2\beta$. In Fig. 14 we plot the quasiparticle densities in the steady state, obtained from the kinetic equation prediction (black dashed line) and MPS simulations (scatter points). We use a maximum MPS bond dimension of $d=100$ and run the simulations starting from the exact ground state, as described in Section III. We fix the noise strength $\gamma_{d}=\gamma_{\phi}=0.1\theta^{2}$, and show results for three different coupling strengths, $\theta=0.2,0.1,0.05$, finding good qualitative agreement for the smaller coupling values $\theta=0.1,0.05$.
Figure 14: Quasiparticle occupation numbers $n^{\infty}_{k}$ in steady state, as a function of quasienergy $\epsilon_{k}$, for noise strength $\gamma_{d}=\gamma_{\phi}=0.1\theta^{2}$. (a) PM and (b) AFM phases. Rate equation prediction (dashed line) vs. MPS simulations (points), for three values of the coupling $\theta$. Protocol parameters are as in Fig. (4). The MPS simulations use a maximum bond dimension of $d=100$.

### v2

@@ D.2 $\sigma^{\pm}_{j;\alpha;\beta}$ matrix elements
We now turn to quasiparticle processes mediated by the $\hat{\sigma}^{\pm}_{j}$ matrix elements. We focus on $\hat{\sigma}^{+}_{j}$ for illustration. In the fermion basis, $\hat{\sigma}^{+}_{j}$ comes attached with the Jordan-Wigner string operator,
$\sigma^{+}_{j;\vec{\alpha};\vec{\beta}}=\bra{\vec{\alpha}}\Big{(}\prod_{i=1}^{j-1}e^{i\pi\hat{c}^{\dagger}_{i}\hat{c}_{i}}\Big{)}\hat{c}_{j}\ket{\vec{\beta}}.$ (105)
Due to this non-locality in the fermionic eigenbasis, multi-particle processes are allowed with $\hat{\sigma}_{j}^{+}$ changing the occupation numbers of up to $12j-$ 1 quasiparticle levels. Amazingly, the exact finite-volume matrix elements have been worked out for the case of the $\hat{X}_{j}$ operators in the TFIM with periodic boundary conditions [89]. Here, we focus on singe- and two-particle transitions near the ground state, which capture the dominant cooling processes at late times, as described in the main text. Our theory can be straightforwardly generalised to include higher-order processes.
@@
$\sigma^{+}_{j;0;k}=\bra{\Omega}\hat{\sigma}^{+}_{j}\ket{k}$ (109)
being the amplitude to directly annihilate a quasiparticle of quasimomentum $k$. We rewrite the single-particle element in fermion basis
$|\sigma^{+}_{j;0;k}|^{2}=|\langle\prod_{i=1}^{j-1}\hat{a}_{2i-1}\hat{a}_{2i}\ \hat{c}_{j}\hat{\eta}^{\dagger}_{k}\rangle_{\Omega}|^{2}.$ (110)
@@
Figure 11: (a) Single-quasiparticle matrix elements $|\sigma^{+}_{j;0;k}|^{2}$ summed over quasimomenta $k$ and plotted against operator position $j$, for different values of $J/g$. (b) Two-particle matrix elements $|\sigma^{+}_{j;0;(k,q)}|^{2}$ summed over $k$, $q$, and plotted against operator position $j$. Note that the two-particle matrix element is non-zero only in the AFM phase.
Two-quasiparticle processes: Based on the physical picture, we expect that in the AFM phase, a pair of domain walls may collide in the bulk and annihilate, as this requires only local reconfigurations of spins. However, there is a subtlety, from the fact that $\hat{\sigma}^{+}_{j}$ has non-zero matrix elements only between eigenstates of different fermion parity. This rules out two-particle transitions in the PM phase, but as discussed in the main text, the issue is sidestepped in the AFM by interacting with the Majorana zero mode, whose action maps between the two nearly-degenerate ground states. We therefore write
$|\sigma^{+}_{j;0;k,q}|^{2}=|\bra{\Omega_{+}}\prod_{i=1}^{j-1}\hat{a}_{2i-1}\hat{a}_{2i}\ \hat{c}_{j}\hat{\eta}^{\dagger}_{k}\hat{\eta}^{\dagger}_{q}\ket{\Omega_{-}}|^{2},$ (111)
with |$\ket{\Omega_{-}}=\hat{\eta}^{\dagger}_{0}\ket{\Omega_{-}}$, and $\hat{\eta}^{\dagger}_{0}$ the Majorana mode. These matrix elements may be evaluated with Wick’s theorem on the $(2j+2)\times(2j+2)$ dimensional $\Sigma$-matrix. In Fig. 11b, we show how the summed matrix elements $\sum_{k,q}|\sigma^{+}_{j;0;k,q}|^{2}$ depend on the position $j$, for different values of $J/g$ in the AFM phase. The pair matrix element is almost homogeneous in the bulk, vanishing in the edge region.
@@
$|\sigma^{+}_{j;k;q}|^{2}=|\bra{\Omega_{+}}\hat{\eta}_{k}\prod_{i=1}^{j-1}\hat{a}_{2i-1}\hat{a}_{2i}\ \hat{c}_{j}\hat{\eta}^{\dagger}_{q}\ket{\Omega_{-}}|^{2}.$ (112)
@@ E.1 Free fermion simulations
The reset of the auxiliaries can be done by reinstating condition (116). If we denote the corresponding map by $\mathcal{R}$, then the cooling channel (1) acts on the covariance matrix according to:
@@ E.2 Tensor network simulations
Here, we summarize the tensor network algorithm that was employed to simulate the dissipative dynamics in Sections III-IV. We give an overview of the relevant details, assuming the reader is familiar with the basics of matrix product algorithms; introductions can be found in e.g. [92, 93].
@@
The time evolution is performed via the standard TEBD algorithm [68, 69], which updates the MPS in real time, followed by the reset channel acting on the bath, which can be efficiently implemented as discussed below. In Fig. 12 we illustrate the evolution for the example of the Floquet TFIM (5) with bath evolution and interaction, Eqs. (7-43). We first simultaneously perform the single-site unitary channels for the system and bath, then the two-body system channels, which correspond to next-nearest-neighbor gates, followed by the nearest-neighbor system-bath channel. Note that the next-nearest-neighbour gate arises from the quasi-1d geometry. This is the most expensive step of the algorithm as it requires two singular value decompositions and scales as $O(\text{dim}(m)^{4}d^{3})$.
Figure 15: Energy of Heisenberg spin ladder vs. rescaled time $t=\nu\theta^{2}$, where $\nu$ is the number of resets and $\theta=0.15$ is the coupling strength. Parameters are the same as in Fig. 6, except here we vary the system size, with ladders of $2\times L$ qubits, with $3\leq L\leq 9$. The data shows that cooling dynamics is independent of system size for the model considered.
@@
Next, we study the effects of truncation due to finite bond dimension during the cooling protocol in Figures 13 and 14. In Fig. 13 we consider the TFIM cooling for the parameters shown in Fig. 3 of the main text. We only show data for two values of the coupling, $\theta=0.1,0.05$ for the MCP, since the SCP protocol is well converged for all bond dimensions. We observe that the convergence is slower for the larger coupling, suggesting higher bond dimensions are needed to capture non-perturbative effects. Phenomenologically, such behavior is expected since higher-order in $\theta$ corrections are expected to lead processes that are less spatially local, thus requiring a larger bond dimension to be accurately captured.
@@
Appendix F System-size dependence for Heisenberg ladder cooling
Here we provide data for system-size dependence of cooling in the Heisenberg spin ladder, discussed in Section IV. We take the same parameters as in Section IV, i.e. $J=J_{\perp}=0.2$, $\theta=0.15$, and $T=\beta=30$. Since the quasiparticle excitations are quasi-local (adiabatically connected to local operators), we expect that both the cooling time and the final energy density scale independently of system size. In Fig. 15 we show how the mean energy (averaged over 60 quantum trajectories) of the $H_{1}$ Floquet Hamiltonian decays, for systems with sizes of $2\times L$ qubits, with $3\leq L\leq 9$. We fix the number of auxiliary qubits at $N_{A}=\text{floor}(L/2)+1$. The data clearly shows that the cooling behaviour is independent of system size, and cooling is achieved in time $O(\theta^{-2})$.
Figure 16: Quasiparticle occupation numbers $n^{\infty}_{k}$ in steady state, as a function of quasienergy $\epsilon_{k}$, for noise strength $\gamma_{d}=\gamma_{\phi}=0.1\theta^{2}$. (a) PM and (b) AFM phases. Rate equation prediction (dashed line) vs. MPS simulations (points), for three values of the coupling $\theta$. Protocol parameters are as in Fig. 4. The MPS simulations use a maximum bond dimension of $d=100$.
Appendix G Comparison of noisy kinetic equation and MPS simulations
In Section V we derived the kinetic equation in the presence of noise, Eq. (64), which includes transitions induced by qubit dephasing and decay. Here we compare the accuracy of the rate equation predictions against MPS simulations. We focus on the MCP with the same parameters as in Fig. 4, namely finite density of auxiliaries, $N_{A}=N_{S}=20$, and $T=\beta=12$. In Fig. 16 we plot the quasiparticle densities in the steady state, obtained from the kinetic equation prediction (black dashed line) and MPS simulations (scatter points). We use a maximum MPS bond dimension of $d=100$ and run the simulations starting from the exact ground state, as described in Section III. We fix the noise strength $\gamma_{d}=\gamma_{\phi}=0.1\theta^{2}$, and show results for three different coupling strengths, $\theta=0.2,0.1,0.05$, finding good qualitative agreement for the smaller coupling values $\theta=0.1,0.05$.